OpenStudy (anonymous):

Improvements at a harley-davidson plant are estimated to be $7.8 million. Construction is expected to take 3 years. What is the future worth of the project in year 3 at an interest rate of 6% per year compounded quarterly, assuming the funds are allocated (a) completely at time 0, and (b) equally at the end of each year?

OpenStudy (anonymous):

This is what i'm wanting to do: Part A: F=7,800,000(F/P,6%,12) = 11,064,300 Part B: A=11,064,300(A/F,6%,12) = 1,586,178.048

OpenStudy (anonymous):

@wolf1728

OpenStudy (wolf1728):

Gee, you started another question? Did you test my other answer?

OpenStudy (anonymous):

Yes, it seemed to make sense. It took me a while to understand it, but i finally did.

OpenStudy (wolf1728):

Well okay - now for this one.

OpenStudy (wolf1728):

$7.8 million 3 years at 6% compounded quarterly Gonna be away from keyboard for a while - taking my dog for a walk. :-)

OpenStudy (wolf1728):

6 % compounded quarterly = 6.136355062% Total = Principal x Rate ^ years Total = 7.8 million * (1.06136355062)^3 At the end of the 3 years Total = 9,325,821.74

OpenStudy (anonymous):

So how would you do part b then?